A store owner purchased 20 halogen light bulbs and 30 fluorescent light bulbs for a total cost of $630. A second purchase, at the same prices, included 30 halogen bulbs and 10 fluorescent bulbs for a total cost of $560. Find the cost of a halogen bulb and a fluorescent bulb.
Answer by jorel1380(3719)

20m+30n=630
30m+10n=560
90m+30n=1680
70m=1050
m=15
n=11
Halogen bulbs cost $15 each. Fluorescent bulbs are $11 each..